A -6 nC particle is moving at 4.0x10^4m/s along the positive z-axis.It passes a region undeflected where there is uniform electric and magnetic fields.The electric field is 5.75 x10^5 N/C along the negative x-axis.What is the magnitude and direction of the magnetic field.


1
Expert's answer
2022-03-22T15:46:19-0400

We know that

"qE=q(V\times B)"

"E=(V\times B)"

"E=VB sin\theta"

"B=\frac{E}{V}"


"B=\frac{5.75\times10^5}{4\times10^4}=14.375T"

Direction of magnetic field

+ y direction